OpenStudy (anonymous):

Find the derivative of the function. f(t) = √(8 – 3t) f '(t) =

OpenStudy (anonymous):

what do you do when the outside function is = √x?

OpenStudy (rogue):

Use the chain rule.\[\frac {d}{dx} u^{n} = n u^{n-1} \frac {du}{dx}\]In your case, \[u = 8 - 3t\]

OpenStudy (anonymous):

i am very confused on how to use the chain rule...

OpenStudy (anonymous):

the notation is killing me. what does the chain rule do in words?

OpenStudy (anonymous):

what trips me up in class is the notation, then i get lost and it al becomes a big jumbled mess